OpenStudy (anonymous):

The greatest of four consecutive even integers is 14 less than twice the smallest integer. What are the intergers?

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

do you understand it

OpenStudy (anonymous):

does any one understand it

OpenStudy (texaschic101):

4 consecutive even integers... x, x + 2, x + 4, x + 6 the greatest is 14 less then twice the smallest.. x + 6 = 2x - 14 6 + 14 = 2x - x 20 = x x + 2 = 20 + 2 = 22 x + 4 = 20 + 4 = 24 x + 6 = 20 + 6 = 26 check.. x + 6 = 2x - 14 20 + 6 = 2(20) - 14 26 = 40 - 14 26 = 26 (correct) so your 4 consecutive even integers are : 20,22,24,26

OpenStudy (anonymous):

Thank you

OpenStudy (texaschic101):

sure thing :)